# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 2.1.3 Parametric amplification
In this section, we introduce another method, which uses parametric amplification to enhance the nonlinear optomechanical interaction Bartkowiak et al. (2014); Lü et al. (2015). Before discussing this method, let us first recall degenerate parametric amplification, which is a very common nonlinear process in quantum optics.
@@
Figure 3: (a) Schematics of using parametric amplification to enhance the optomechanical interaction. A cavity, with frequency $\omega_{\rm cav}$, contains a $\chi^{\left(2\right)}$ nonlinear crystal, which is driven by a two-photon driving of frequency $\omega_{\rm 2ph}$, amplitude $\Omega_{\rm 2ph}$, and phase $\theta_{\rm 2ph}$. The right mirror of the cavity is oscillating with frequency $\omega_{m}$ and position $x\left(t\right)$. (b) Enhanced optomechanical interaction $g_{\rm sq0}$ (left axis) and resonance frequency of the squeezed cavity mode $\omega_{\rm sq}$ (right axis) as a function of the detuning $\Delta_{\rm 2ph}$ for $\Omega_{\rm 2ph}=4000\omega_{m}$. The inset shows the ratios $g_{\rm sq0}/\omega_{m}$ and $g_{\rm sq0}^{\prime}/\omega_{\rm sq}$. Panels (a) and (b) are reproduced with permission from Ref. Lü et al. (2015), X.-Y. Lü et al., Phys. Rev. Lett. 114, 093602 (2015).
The nonlinear transformation in Eq. (38) shows a Bogoliubov coefficient of $\cosh\left(r\right)$, which increases exponentially with the squeezing parameter $r$. This indicates that in the squeezed frame, the coupling of the cavity mode to other degrees of freedom can be enhanced exponentially. Such a mechanism has been used for enhancing the optomechanical interaction in cavity-optomechanical systems Lü et al. (2015). A schematic setup is illustrated in Fig. 3(a). Crucially, a nonlinear $\chi^{\left(2\right)}$ crystal is placed inside a cavity, such that the cavity mode is subject to a detuned two-photon driving. As mentioned above, the cavity mode $a$ is squeezed by the two-photon driving and, as a result, becomes the squeezed mode $a_{\rm sq}$. Expressed in terms of the mode $a_{\rm sq}$, the optomechanical interaction Hamiltonian in Eq. (5) is transformed to
@@
This implies that, compared to the single-photon coupling $g_{0}$ in the original laboratory frame, an exponential enhancement of $g_{\rm sq}^{\rm om}$ can be achieved with increasing the squeezing parameter $r$, i.e., when $\Delta_{\rm 2ph}$ approaches $\Omega_{\rm 2ph}$ from the right, as seen in Fig. 3(b). Under a similar mechanism, the quadratic optomechanical interaction can also be exponentially enhanced Gu et al. (2018); Zhang et al. (2019).
@@
while $\gamma_{m}$ is the mechanical decay rate and $n_{\rm th}$ is the thermal phonon number of the mechanical mode. It is seen that by increasing the squeezing parameter $r$, the optomechanical system can be driven into the single-photon strong coupling regime (i.e., $g_{\rm sq}^{\rm om}>\kappa$) even from the single-photon WC regime ($g_{0}<\kappa$). As a direct result, such a parametric enhancement of the optomechanical interaction improves the conversion Luo et al. (2020), entanglement Xiong et al. (2017), and even cross-Kerr nonlinearity Yin et al. (2018) between the optical and microwave fields. The two-mode squeezing of the cavity field is also capable of exponentially enhancing the optomechanical interaction, but at the same time a two-mode, rather than single-mode, squeezed vacuum field is needed to suppress the noise induced by this two-mode squeezing Li et al. (2016).
@@
which can be used to simulate the dynamical Casimir effect Qin et al. (2019); Wilson et al. (2011). In fact, the dynamics described by $H_{-}^{\rm sq}$ can be interpreted in the laboratory frame as mechanically induced two-photon hyper-Raman scattering, i.e., an anti-Stokes scattering of a driving photon pair, rather than a single photon, into a higher-energy mode by absorbing a phonon.
Furthermore, by squeezing the cavity field, one can also eliminate the quantum backaction heating even in the unresolved sideband regime, such that ground-state cooling Lau and Clerk (2020) and mechanical squeezing Xiong et al. (2020) can be implemented. Instead, squeezing the mechanical mode provides another way to enhance the optomechanical interaction. For examples, along this line, photon blockade Lemonde et al. (2016); Yin et al. (2017), mechanical squeezing Zhang et al. (2018b), superradiant quantum phase transition Lü et al. (2018), and enhancing quadratic optomechanical interaction Gu et al. (2018) have been studied.
Recently, Ref. Arenz et al. (2020) showed that by combining parametric amplification processes and dynamical-decoupling techniques, one can amplify the desired interaction, and at the same time suppress an undesired interaction, so as to speed up the dynamical evolution of the system.
@@ 2.1.4 Other amplification mechanisms
In contrast to conventional Hermitian Hamiltonians, non-Hermitian parity-time ($\mathcal{PT}$) symmetric Hamiltonians Bender and Boettcher (1998); El-Ganainy et al. (2018); Özdemir et al. (2019) exhibit a phase transition from the unbroken to broken $\mathcal{PT}$ phases at an exceptional point, where the eigenvalues are changed from real to complex numbers. By coupling two different optical or microwave cavities, one with passive loss and the other with active gain, $\mathcal{PT}$-symmetric systems can be created Peng et al. (2014); Hodaei et al. (2014); Chang et al. (2014). In such double-cavity systems, many counterintuitive aspects occur, particularly in the vicinity of exceptional points, which can be explored to implement a strong nonlinearity for cavity optomechanics. By manipulating the gain-to-loss ratio, a nonlinear regime for the intracavity-photon intensity can emerge, such that the optical pressure and the mechanical gain are enhanced simultaneously, resulting in an ultralow-threshold phonon laser Jing et al. (2014); Zhang et al. (2018a). A $\mathcal{PT}$-symmetry-induced enhancement mechanism has also been used for demonstrating optomechanical chaos Lü et al. (2015), optomechanically-induced transparency Jing et al. (2015), and high-precision metrology Liu et al. (2016).
@@
By coupling a cavity field to an array of mechanical resonators, a collectively enhanced optomechanical interaction can be obtained Gärtner et al. (2018); Piergentili et al. (2018). It has been predicted in Refs. Xuereb et al. (2012, 2013) that the single-photon coupling strength of the cavity mode to a collective mechanical mode can scale as $N^{3/2}$, where $N$ is the number of mechanical resonators. Compared to the case of a single mechanical resonator, the resulting collective coupling can be made up to several orders of magnitude stronger, which, as a direct consequence, enables exploiting the long-range interactions between distant mechanical resonators. Optimal configurations for these optomechanical arrays are given in Ref. Newsom et al. (2020), such that the optomechanical interaction strength scales exponentially with $N$ before the saturation. Furthermore, the collective enhancement of the single-photon optomechanical interaction has also been shown in the cases where the collective motion of ensembles of ultracold atoms, serving as a mechanical resonator, is coupled to the cavity field Brennecke et al. (2008); Murch et al. (2008); Brooks et al. (2012); Ritsch et al. (2013).
@@
For microwave-regime superconducting cavities, the mechanical motion, which modulates the capacitance of a cavity, can couple to the cavity mode via the radiation-pressure interaction Regal et al. (2008); Teufel et al. (2011); Massel et al. (2011). When a Cooper-pair transistor is embedded inside a superconducting cavity Heikkilä et al. (2014), the single-photon optomechanical interaction strength can be enhanced by several orders of magnitude. Such a giant enhancement arises due to the presence of the nonlinearity of the Josephson junctions, and has been experimentally demonstrated in Ref. Pirkkalainen et al. (2015b). Similarly, the nonlinearity of Coulomb blockade can also be used to significantly enhance the single-photon optomechanical interaction Blien et al. (2020). It has been suggested that, with a Cooper-pair transistor Rimberg et al. (2014) or a Cooper-pair box Manninen et al. (2022), even the USC between cavity photons and a mechanical resonator can be reached. Moreover, an ultrastrong single-photon optomechanical interaction can also be obtained by embedding a dc superconducting quantum interference device (SQUID), with a suspended arm as a mechanical resonator, into a superconducting cavity Shevchuk et al. (2017); Nation et al. (2016); Kounalakis et al. (2020).
@@ 2.2 Amplified photon-atom interactions in cavity quantum electrodynamics
where $g=-\boldsymbol{d}_{ge}\cdot\boldsymbol{\epsilon}=-\boldsymbol{d}_{eg}\cdot\boldsymbol{\epsilon}$ is the coupling strength, while $\sigma^{-}=\left|{g}\rangle\langle{e}\right|$ and $\sigma^{+}=\left|{e}\rangle\langle{g}\right|$ are the ladder operators of the atom. Equation (53) can describe the atom-field interaction, in particular, when the coupling strength $g$ is comparable to the atomic transition frequency or the cavity frequency. The quantum Rabi model is discussed in the following sections. Here, we focus on the case when $g$ is much smaller than these two characteristic frequencies. In this case, one can apply the RWA so as to neglect the counter-rotating components, yielding the JC interaction Hamiltonian Jaynes and Cummings (1963), i.e.,
@@
which shows that the ability, e.g., to process quantum information, increases as the coupling strength $g$. The first demonstration of the strong coupling with single atoms was reported in Ref. Meschede et al. (1985). Below we review several methods for amplifying the coupling strength $g$ and the cooperativity $C$.
@@ 2.2.1 Parametric amplification
The coupling in Eq. (54) essentially originates from the fluctuations of the electromagnetic vacuum Haroche and Raimond (2006), and thus amplifying these fluctuations via antisqueezing could induce an enhancement of the coupling strength $g$. The basic idea underlying this amplification method, essentially the same with the idea presented in Sec. 2.1.3, is shown schematically in Fig. 4(a).
Figure 4: (a) Schematics of using parametric amplification to enhance the photon-atom interaction $g$ and thus the cooperativity $C$. The single-mode cavity contains a $\chi^{\left(2\right)}$ nonlinear medium, which is strongly pumped at amplitude $\Omega_{\rm 2ph}$, frequency $\omega_{\rm 2ph}$, and phase $\theta_{\rm 2ph}$. A two-level atom, with the ground state $\left|{g}\right\rangle$ and the excited state $\left|{e}\right\rangle$, is coupled to the cavity mode with a strength $g$; $\omega_{\rm cav}$ and $\omega_{q}$ are the resonance frequencies of the cavity and the atom, respectively, $\kappa$ is the cavity decay rate, and $\gamma$ is the atomic spontaneous emission rate. (b) Exponentially enhanced cooperativity $C_{\rm sq}$ and coupling $g_{\rm sq}^{\rm rw}$ (inset) between the squeezed cavity mode and the atom. The gray and yellow shaded areas refer to the weak ($C_{\rm sq}<1$) and strong ($C_{\rm sq}>1$) coupling regimes under the assumption of $C=0.2$. (c) Exponentially enhanced entanglement of two three-level $\Lambda$-type atoms mediated by the squeezed cavity mode. With squeezing, the lower bound of the entanglement infidelity $\delta=1-F$ is lowered to $\approx 6/\left(e^{2r}C\right)$, clearly far below $1/C$ and $1/\sqrt{C}$, which are the lower bounds imposed, respectively, by unitary and dissipative state preparation without squeezing. Curves show effective predictions and symbols indicate exact results. Panels (a)-(c) are adapted with permission from Ref. Qin et al. (2018), W. Qin et al., Phys. Rev. Lett. 120, 093601 (2018).
When the cavity field is parametrically driven (i.e., is squeezed), as described by the Hamiltonian in Eq. (35), the coupling strength $g$ can be exponentially enhanced Qin et al. (2018); Leroux et al. (2018); Zeytinoğlu et al. (2017). Upon substituting the Bogoliubov transformation in Eq. (38) into Eq. (54), the atom-cavity coupling Hamiltonian $H_{\rm JC}^{\rm int}$ is transformed to
@@
When entangled states of separated atoms are prepared in optical or microwave cavities, the state infidelity $\delta=1-F$ scales as $\delta\propto 1/\sqrt{C}$ for the preparation approaches based on unitary gates, and as $\propto 1/C$ for dissipative state preparation. Here, $F$ is the fidelity between the actual and ideal states. Thus, the cooperativity enhancement given in Eq. (64), when applied to dissipative state preparation, can enable an exponential improvement in the state infidelity Qin et al. (2018), i.e.,
@@
So far, parametrically amplified photon-atom interactions, and thus mediated atom-atom interactions, have been widely studied. This mechanism has been explored to generate lasing into a squeezed cavity mode Sánchez Muñoz and Jaksch (2021). It has been also shown that when adiabatically eliminating the degree of freedom, an exponential enhancement of the dipole-dipole coupling between atoms can be observed Wang et al. (2019a). If the amplitude of the atom driving is modulated in time, then a fast and high-fidelity generation of steady-state entanglement Chen et al. (2019), and similarly, a high-fidelity implementation of arbitrary phase gates Wang et al. (2020a) can be achieved by a parametric driving of the cavity. In the case of two coupled cavities, it is possible to parametrically drive a single cavity, which is coupled to an atom, to enhance the coupling of this atom to another cavity Wang et al. (2019b). When, furthermore, a multiphoton coupling between the atom and the cavity is taken into account, their single-photon or two-photon coupling can be greatly enhanced even for a small squeezing parameter Wang et al. (2020b). Enhanced spin-phonon, and in turn spin-spin, interactions via squeezing a mechanical resonator have been theoretically demonstrated in hybrid systems in Ref. Li et al. (2020). Very recently, a theoretical proposal that can amplify magnon-spin interactions via virtually-excited squeezed phonons has also been put forward in magnonics Wang et al. (2023); Nori (2023).

### v2

@@ 2.1.3 Parametric amplification
In this section, we introduce another method, which uses parametric amplification to enhance the nonlinear optomechanical interaction [210, 246]. Before discussing this method, let us first recall degenerate parametric amplification, which is a very common nonlinear process in quantum optics.
@@
Figure 3: (a) Schematics of using parametric amplification to enhance the optomechanical interaction. A cavity, with frequency $\omega_{\rm cav}$, contains a $\chi^{\left(2\right)}$ nonlinear crystal, which is driven by a two-photon driving of frequency $\omega_{\rm 2ph}$, amplitude $\Omega_{\rm 2ph}$, and phase $\theta_{\rm 2ph}$. The right mirror of the cavity is oscillating with frequency $\omega_{m}$ and position $x\left(t\right)$. (b) Enhanced optomechanical interaction $g_{\rm sq0}$ (left axis) and resonance frequency of the squeezed cavity mode $\omega_{\rm sq}$ (right axis) as a function of the detuning $\Delta_{\rm 2ph}$ for $\Omega_{\rm 2ph}=4000\omega_{m}$. The inset shows the ratios $g_{\rm sq0}/\omega_{m}$ and $g_{\rm sq0}^{\prime}/\omega_{\rm sq}$. Panels (a) and (b) are reproduced with permission from Ref. [246], X.-Y. Lü et al., Phys. Rev. Lett. 114, 093602 (2015).
The nonlinear transformation in Eq. (38) shows a Bogoliubov coefficient of $\cosh\left(r\right)$, which increases exponentially with the squeezing parameter $r$. This indicates that in the squeezed frame, the coupling of the cavity mode to other degrees of freedom can be enhanced exponentially. Such a mechanism has been used for enhancing the optomechanical interaction in cavity-optomechanical systems [246]. A schematic setup is illustrated in Fig. 3(a). Crucially, a nonlinear $\chi^{\left(2\right)}$ crystal is placed inside a cavity, such that the cavity mode is subject to a detuned two-photon driving. As mentioned above, the cavity mode $a$ is squeezed by the two-photon driving and, as a result, becomes the squeezed mode $a_{\rm sq}$. Expressed in terms of the mode $a_{\rm sq}$, the optomechanical interaction Hamiltonian in Eq. (5) is transformed to
@@
This implies that, compared to the single-photon coupling $g_{0}$ in the original laboratory frame, an exponential enhancement of $g_{\rm sq}^{\rm om}$ can be achieved with increasing the squeezing parameter $r$, i.e., when $\Delta_{\rm 2ph}$ approaches $\Omega_{\rm 2ph}$ from the right, as seen in Fig. 3(b). Under a similar mechanism, the quadratic optomechanical interaction can also be exponentially enhanced [247, 248].
@@
while $\gamma_{m}$ is the mechanical decay rate and $n_{\rm th}$ is the thermal phonon number of the mechanical mode. It is seen that by increasing the squeezing parameter $r$, the optomechanical system can be driven into the single-photon strong coupling regime (i.e., $g_{\rm sq}^{\rm om}>\kappa$) even from the single-photon WC regime ($g_{0}<\kappa$). As a direct result, such a parametric enhancement of the optomechanical interaction improves the conversion [249], entanglement [250], and even cross-Kerr nonlinearity [251] between the optical and microwave fields. The two-mode squeezing of the cavity field is also capable of exponentially enhancing the optomechanical interaction, but at the same time a two-mode, rather than single-mode, squeezed vacuum field is needed to suppress the noise induced by this two-mode squeezing [252].
@@
which can be used to simulate the dynamical Casimir effect [253, 46]. In fact, the dynamics described by $H_{-}^{\rm sq}$ can be interpreted in the laboratory frame as mechanically induced two-photon hyper-Raman scattering, i.e., an anti-Stokes scattering of a driving photon pair, rather than a single photon, into a higher-energy mode by absorbing a phonon.
Furthermore, by squeezing the cavity field, one can also eliminate the quantum backaction heating even in the unresolved sideband regime, such that ground-state cooling [254] and mechanical squeezing [255] can be implemented. Instead, squeezing the mechanical mode provides another way to enhance the optomechanical interaction. For examples, along this line, photon blockade [256, 257], mechanical squeezing [258], superradiant quantum phase transition [259], and enhancing quadratic optomechanical interaction [247] have been studied.
Recently, Ref. [260] showed that by combining parametric amplification processes and dynamical-decoupling techniques, one can amplify the desired interaction, and at the same time suppress an undesired interaction, so as to speed up the dynamical evolution of the system.
@@ 2.1.4 Other amplification mechanisms
In contrast to conventional Hermitian Hamiltonians, non-Hermitian parity-time ($\mathcal{PT}$) symmetric Hamiltonians [261, 262, 263] exhibit a phase transition from the unbroken to broken $\mathcal{PT}$ phases at an exceptional point, where the eigenvalues are changed from real to complex numbers. By coupling two different optical or microwave cavities, one with passive loss and the other with active gain, $\mathcal{PT}$-symmetric systems can be created [264, 265, 266]. In such double-cavity systems, many counterintuitive aspects occur, particularly in the vicinity of exceptional points, which can be explored to implement a strong nonlinearity for cavity optomechanics. By manipulating the gain-to-loss ratio, a nonlinear regime for the intracavity-photon intensity can emerge, such that the optical pressure and the mechanical gain are enhanced simultaneously, resulting in an ultralow-threshold phonon laser [236, 237]. A $\mathcal{PT}$-symmetry-induced enhancement mechanism has also been used for demonstrating optomechanical chaos [267], optomechanically-induced transparency [268], and high-precision metrology [269].
@@
By coupling a cavity field to an array of mechanical resonators, a collectively enhanced optomechanical interaction can be obtained [270, 271]. It has been predicted in Refs. [272, 273] that the single-photon coupling strength of the cavity mode to a collective mechanical mode can scale as $N^{3/2}$, where $N$ is the number of mechanical resonators. Compared to the case of a single mechanical resonator, the resulting collective coupling can be made up to several orders of magnitude stronger, which, as a direct consequence, enables exploiting the long-range interactions between distant mechanical resonators. Optimal configurations for these optomechanical arrays are given in Ref. [274], such that the optomechanical interaction strength scales exponentially with $N$ before the saturation. Furthermore, the collective enhancement of the single-photon optomechanical interaction has also been shown in the cases where the collective motion of ensembles of ultracold atoms, serving as a mechanical resonator, is coupled to the cavity field [275, 276, 277, 278].
@@
For microwave-regime superconducting cavities, the mechanical motion, which modulates the capacitance of a cavity, can couple to the cavity mode via the radiation-pressure interaction [279, 220, 226]. When a Cooper-pair transistor is embedded inside a superconducting cavity [280], the single-photon optomechanical interaction strength can be enhanced by several orders of magnitude. Such a giant enhancement arises due to the presence of the nonlinearity of the Josephson junctions, and has been experimentally demonstrated in Ref. [281]. Similarly, the nonlinearity of Coulomb blockade can also be used to significantly enhance the single-photon optomechanical interaction [282]. It has been suggested that, with a Cooper-pair transistor [283] or a Cooper-pair box [284], even the USC between cavity photons and a mechanical resonator can be reached. Moreover, an ultrastrong single-photon optomechanical interaction can also be obtained by embedding a dc superconducting quantum interference device (SQUID), with a suspended arm as a mechanical resonator, into a superconducting cavity [285, 286, 287].
@@ 2.2 Amplified photon-atom interactions in cavity quantum electrodynamics
where $g=-\boldsymbol{d}_{ge}\cdot\boldsymbol{\epsilon}=-\boldsymbol{d}_{eg}\cdot\boldsymbol{\epsilon}$ is the coupling strength, while $\sigma^{-}=\left|{g}\rangle\langle{e}\right|$ and $\sigma^{+}=\left|{e}\rangle\langle{g}\right|$ are the ladder operators of the atom. Equation (53) can describe the atom-field interaction, in particular, when the coupling strength $g$ is comparable to the atomic transition frequency or the cavity frequency. The quantum Rabi model is discussed in the following sections. Here, we focus on the case when $g$ is much smaller than these two characteristic frequencies. In this case, one can apply the RWA so as to neglect the counter-rotating components, yielding the JC interaction Hamiltonian [28], i.e.,
@@
which shows that the ability, e.g., to process quantum information, increases as the coupling strength $g$. The first demonstration of the strong coupling with single atoms was reported in Ref. [288]. Below we review several methods for amplifying the coupling strength $g$ and the cooperativity $C$.
@@ 2.2.1 Parametric amplification
The coupling in Eq. (54) essentially originates from the fluctuations of the electromagnetic vacuum [5], and thus amplifying these fluctuations via antisqueezing could induce an enhancement of the coupling strength $g$. The basic idea underlying this amplification method, essentially the same with the idea presented in Sec. 2.1.3, is shown schematically in Fig. 4(a).
Figure 4: (a) Schematics of using parametric amplification to enhance the photon-atom interaction $g$ and thus the cooperativity $C$. The single-mode cavity contains a $\chi^{\left(2\right)}$ nonlinear medium, which is strongly pumped at amplitude $\Omega_{\rm 2ph}$, frequency $\omega_{\rm 2ph}$, and phase $\theta_{\rm 2ph}$. A two-level atom, with the ground state $\left|{g}\right\rangle$ and the excited state $\left|{e}\right\rangle$, is coupled to the cavity mode with a strength $g$; $\omega_{\rm cav}$ and $\omega_{q}$ are the resonance frequencies of the cavity and the atom, respectively, $\kappa$ is the cavity decay rate, and $\gamma$ is the atomic spontaneous emission rate. (b) Exponentially enhanced cooperativity $C_{\rm sq}$ and coupling $g_{\rm sq}^{\rm rw}$ (inset) between the squeezed cavity mode and the atom. The gray and yellow shaded areas refer to the weak ($C_{\rm sq}<1$) and strong ($C_{\rm sq}>1$) coupling regimes under the assumption of $C=0.2$. (c) Exponentially enhanced entanglement of two three-level $\Lambda$-type atoms mediated by the squeezed cavity mode. With squeezing, the lower bound of the entanglement error $\delta=1-F$ is lowered to $\approx 6/\left(e^{2r}C\right)$, clearly far below $1/C$ and $1/\sqrt{C}$, which are the lower bounds imposed, respectively, by unitary and dissipative state preparation without squeezing. Curves show effective predictions and symbols indicate exact results. Panels (a)-(c) are adapted with permission from Ref. [289], W. Qin et al., Phys. Rev. Lett. 120, 093601 (2018).
When the cavity field is parametrically driven (i.e., is squeezed), as described by the Hamiltonian in Eq. (35), the coupling strength $g$ can be exponentially enhanced [289, 290, 291]. Upon substituting the Bogoliubov transformation in Eq. (38) into Eq. (54), the atom-cavity coupling Hamiltonian $H_{\rm JC}^{\rm int}$ is transformed to
@@
When entangled states of separated atoms are prepared in optical or microwave cavities, the state error $\delta=1-F$ scales as $\delta\propto 1/\sqrt{C}$ for the preparation approaches based on unitary gates, and as $\propto 1/C$ for dissipative state preparation. Here, $F$ is the fidelity between the actual and ideal states. Thus, the cooperativity enhancement given in Eq. (64), when applied to dissipative state preparation, can enable an exponential improvement in the state error [289], i.e.,
@@
So far, parametrically amplified photon-atom interactions, and thus mediated atom-atom interactions, have been widely studied. This mechanism has been explored to generate lasing into a squeezed cavity mode [292]. It has been also shown that when adiabatically eliminating the degree of freedom, an exponential enhancement of the dipole-dipole coupling between atoms can be observed [293]. If the amplitude of the atom driving is modulated in time, then a fast and high-fidelity generation of steady-state entanglement [294], and similarly, a high-fidelity implementation of arbitrary phase gates [295] can be achieved by a parametric driving of the cavity. In the case of two coupled cavities, it is possible to parametrically drive a single cavity, which is coupled to an atom, to enhance the coupling of this atom to another cavity [296]. When, furthermore, a multiphoton coupling between the atom and the cavity is taken into account, their single-photon or two-photon coupling can be greatly enhanced even for a small squeezing parameter [297]. Enhanced spin-phonon, and in turn spin-spin, interactions via squeezing a mechanical resonator have been theoretically demonstrated in hybrid systems in Ref. [298]. Very recently, a theoretical proposal that can amplify magnon-spin interactions via virtually-excited squeezed phonons has also been put forward in magnonics [211, 212].